What is probability of getting head and tail alternatively in three throws of a coin?
If you mean 1: head 2:tail 3:head the probability is 1/2 x 1/2 x 1/2 = 1/8.
If you mean only that you want the heads and tails to alternate, that is, you mean the above or tail, head , tail, the probability is 1/8 + 1/8 = 1/4.
What is the probability of getting male and female offsring in human?
The global gender probability, at birth, is approx 0.517 for male and 0.483 for female. This is not the same as the gender ratio at conception because of gender-specific abortion.
What are the possible outcomes if three coins are tossed and die is rolled?
There are 48 possible outcomes and I have no intention of listing them all.
They are all of the form CCCD
where C = H or T, and D takes the numeric values from 1 to 6.
What is the difference between standard deviation and mean?
The mean is the average value and the standard deviation is the variation from the mean value.
64/425, or approximately .151
P(exactly 3 hearts)=(13/52)*(12/51)*(11/50)=11/850
P(exactly 2 hearts)=(13/52)*(12/51)*(39/50)+(13/52)*(39/51)*(12/50)+(39/52)*(13/51)*(12/50)=3*(13/52)*(12/51)*(39/50)=117/850
P(at least 2 hearts)=P(exactly 3 hearts)+P(exactly 2 hearts)=11/850+117/850=128/850=64/425, or approximately .151
If you draw 5 cards from a standard deck of cards what is the probability of drawing 5 hearts?
Probability (P)of drawing 5 hearts (H) is: P(H) and P(H) and P(H) and P(H) and P(H). There are 13 hearts in a 52 card deck. So, probability is: 13/52 * 12/51 * 11/50 * 10/49 * 9/48 or 154440/311875200 or 0.000495.
The expected (but incorrect) answer to the question is 1/2.
It is not correct because:
What is the difference between probability and membership function?
Membership is Not the same as probability.v Probability describes the uncertainty of event occurrence .The probability is an uncertainty associated with time.
vFuzziness describes event ambiguity , i.e., the degree to which an event occurs not whether it occurs
What is the probability that three coin flips will have at least one head?
Pr(3 flips at least one H) = 1 - Pr(3 flips, NO heads)
= 1 - Pr(3 flips, TTT) = 1 - (1/2)3 = 1 - 1/8 = 7/8
What is the probability of getting two 1's when two number cubes are rolled?
The probability is 1/36.
Is probability density function and gaussian distribution function are same?
yes, h=1/sigma(standard deviation)
How do you write 21 as a probability?
21 is a number. You cannot write that as a probability since it is not an event.
Is drawing two cards from a regular deck of cards and not replacing them dependent or independent?
dependent
What is the probability of rolling an odd number less than 5?
If you're rolling an honest, normal cube, labeled 1 through 6 . . .
-- There are 6 possible outcomes of one roll.
-- Two of them (1 and 3) are odd numbers less than 5.
The probability is 2/6 = 1/3 = 331/3%.
The 'odds' are 2 to 1 against it.
What are some limitations of statistics?
limitations of statistics are as follows:
1. Statistics does not deal with an individual
2.It is not suitable to the study of qualitative phenomenon
3.Statistical relations are not exact
4.Statistics is liable to be misused
5.Statistics is only a means
How many people have the same birthday in the world?
It must be about 360,000 - the approximate number of world births per day
it is estematet to be that 150 million or billion(i cant remember) people share your birthday
Is grouped data better than ungrouped data?
Groued or ungrouped data all used to conclude something,but that conclusion is not 100% accurate,Whatever we use those just to live in this world,so dont bother about it . Groued or ungrouped data all used to conclude something,but that conclusion is not 100% accurate,Whatever we use those just to live in this world,so dont bother about it .
From the cross AA x AA the probability of producing a homozygous dominant offspring is?
From the cross Aa x Aa, the probability of producing a homozygous dominant offspring is 1/4 or (0.25).
How do you interpret Probability of default?
Distance-to-Default:-
- distance between the expected value of the asset and the default point
- after substitution into a normal c.d.f one gets probability of default
DD(t) =ln(V/F)+(µ-0.5*σ^2)*T/(σ*sqrt(T));
Where, V= value of the assets
F=Value of the liability/debt
µ= expected return of assets
σ=Volatility of the assets
T= Time
And Probability of default:-
PD(t) = NormDist(-DD)= Ɲ(-DD)
What is the formula to convert normal distribution to the standard normal distribution?
If a variable X, is distributed Normally with mean m and standard deviation s then
Z = (X - m)/s has a standard normal distribution.
What is a probability that a child would be a girl versus a boy?
It would depend on the parents' genes and age. Globally, the probability is approx 0.483